E2-19 Calculating Missing Amounts and Cost of Goods Manufactured and Sold [LO 2-3, 2-4, 2-5, 2-6] For each of the following independent cases (1-4), compute the missing values. (Enter all amounts as positive values.) Case 1 Case 2 Case 3 Case 4 $6,400 55,000 700 1,600 $ 22,000$91,000 Beginning raw materials Raw material purchases Indirect materials issued Ending raw materials Direct materials used Direct labor Manufacturing overhead Total current manufacturing costs Beginning work in process Ending work in process Cost of goods manufactured Beginning finished goods Ending finished goods Cost of goods sold 11,350 1,100 2,350 14,400 41,640 600 2,400 95,200 32,220 37,000 122,950 39,950 81,600 36.600 59,000 30,280 540,030 93,100 42,000 41,350 89,725 53,000 102,220 237,900 825,900 135,000 76,000 72,000 73,800 31,200 398,500 73,000 111,775 842,400
Answer #1

Case 1

Case 2

Case 3

Case 4

Beginning raw material

$        6,400.00

$    6,500.00

$      22,000.00

$      91,000.00

Raw material purchases

$      55,000.00

$ 11,350.00

$      41,640.00

$ 2,96,200.00

Indirect material issued

$            700.00

$    1,100.00

$            600.00

$        2,400.00

Ending raw material

$        1,600.00

$    2,350.00

$      30,820.00

$      95,200.00

Direct material used

$      59,100.00

$ 14,400.00

$      32,220.00

$ 2,98,600.00

Direct labor

$      37,000.00

$ 27,250.00

$      30,600.00

$ 1,22,950.00

Manufacturing overheads

$      59,000.00

$ 39,950.00

$      30,280.00

$ 5,40,030.00

Total Current manufacturing costs

$ 1,55,100.00

$ 81,600.00

$      93,100.00

$ 9,61,580.00

Beginning Work in process

$      42,000.00

$ 36,600.00

$      37,975.00

$ 1,02,220.00

Ending Work in process

$      62,100.00

$ 44,400.00

$      41,350.00

$ 2,37,900.00

Cost of goods manufactured

$ 1,35,000.00

$ 73,800.00

$      89,725.00

$ 8,25,900.00

Beginning finished goods

$      76,000.00

$ 30,400.00

$      53,000.00

$ 4,06,000.00

Ending finished goods

$      72,000.00

$ 31,200.00

$      30,950.00

$ 3,89,500.00

Cost of goods sold

$ 1,39,000.00

$ 73,000.00

$ 1,11,775.00

$ 8,42,400.00

Working note

Case 1

Case 2

Case 3

Case 4

Beginning raw material

6400

=14400-11350+1100+2350

22000

91000

Raw material purchases

55000

11350

41640

=2400+95200+289600-91000

Indirect material issued

700

1100

600

2400

Ending raw material

1600

2350

=22000+41640-32220-600

95200

Direct material used

=6400+55000-700-1600

14400

32220

=961580-540030-122950

Direct labor

37000

=81600-14400-39950

=93100-32220-30280

122950

Manufacturing overheads

59000

39950

30280

540030

Total Current manufacturing costs

=59100+37000+59000

81600

93100

=825900-102220+237900

Beginning Work in process

42000

36600

=89725+41350-93100

102220

Ending Work in process

=155100+42000-135000

=81600+36600-73800

41350

237900

Cost of goods manufactured

135000

73800

89725

825900

Beginning finished goods

76000

=73000+31200-73800

53000

=842400+389500-825900

Ending finished goods

72000

31200

=89725+53000-111775

389500

Cost of goods sold

=135000+76000-72000

73000

111775

842400